A chemist needs 130 milliliters of a 47% solution but has only 35% and 87% solutions available. How many milliliters of each should be mixed to get the desired solution?
Answer by Alan3354(69443)

A chemist needs 130 milliliters of a 47% solution but has only 35% and 87% solutions available. How many milliliters of each should be mixed to get the desired solution?
=============
t = amount of 35%
e = amount of 87&
---
t + e = 130 (total solution)
35t + 87e = 47*130 (total soluent)
========================
Solve for t & e